# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v3

@@ Abstract
We study stopping rules for stochastic gradient descent (SGD) for convex optimization from the perspective of anytime-valid confidence sequences. Classical analyses of SGD provide convergence guarantees in expectation or at a fixed horizon, but offer no statistically valid way to assess, at an arbitrary time, how close the current iterate is to the optimum. We develop an anytime-valid, data-dependent upper confidence sequence for the weighted average suboptimality of projected SGD, constructed via nonnegative supermartingales and requiring no smoothness or strong convexity. This confidence sequence yields a simple stopping rule that is provably $\varepsilon$-optimal with probability at least $1-\alpha$, with explicit bounds on the stopping time under standard stochastic approximation stepsizes. To the best of our knowledge, these are the first rigorous, time-uniform performance guarantees and finite-time $\varepsilon$-optimality certificates for projected SGD with general convex objectives, based solely on observable trajectory quantities.
@@ 1 Introduction
Stochastic gradient methods are the workhorses of modern optimization and machine learning. Their appeal lies in a rare combination of simplicity, scalability, and robustness to noisy information. Yet, despite decades of research, a persistent gap remains between theory and practice: during an optimization run, practitioners have no data-dependent, statistically valid indication of how close the current iterate is to optimality.
Classical analyses provide valuable insights into the behavior of stochastic gradient descent (SGD), including convergence in expectation, asymptotic rates, and fixed-horizon high-probability bounds. However, these guarantees typically rely on problem-dependent constants, such as smoothness parameters, strong convexity coefficients, and variance bounds, which are unknown in practice, and they offer no protection under adaptive monitoring or early stopping. Thus, when training a model or tuning a hyperparameter, a user is left with the fundamental but unanswered question:
How close am I to the optimum right now?
In practical settings, such as large-scale training, reinforcement learning, scientific computing, or any computation where time or data is limited, this question is more than academic. Stopping too early risks under-optimization, while stopping too late wastes considerable resources. Yet most stopping rules used in practice are heuristic: they monitor validation curves, impose fixed budgets, or rely on experience rather than statistical guarantees.
In this paper, we revisit this classical challenge from a sequential-inference perspective. The key idea is to view the progress of SGD as a stochastic process that can be monitored through anytime-valid confidence sequences (CSs). A confidence sequence is a sequence of intervals that remains valid uniformly over time, even under continuous monitoring and adaptive stopping. By adapting this framework to stochastic optimization, we obtain fully data-dependent, time-uniform upper bounds on the suboptimality gap, requiring neither knowledge of the optimization horizon nor precise knowledge of problem-dependent parameters. These guarantees immediately yield statistically valid, data-dependent stopping rules that certify $\varepsilon$-optimality under arbitrary adaptive monitoring.
@@ 1.1 Problem formulation
To formalize the problem, we consider optimization over a closed and convex feasible region. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a nonempty compact convex set, and let $f:\mathcal{X}\to\mathbb{R}$ be a convex objective function. We denote an arbitrary minimizer of $f$ over $\mathcal{X}$ by
$x^{\star}\in\arg\min_{x\in\mathcal{X}}f(x).$
We study the classical projected variant of stochastic gradient descent, defined recursively by
$x_{t+1}=\Pi_{\mathcal{X}}(x_{t}-\eta_{t}g_{t}),$ (1)
where $\Pi_{\mathcal{X}}$ denotes Euclidean projection onto $\mathcal{X}$ and $\{\eta_{t}\}_{t\geq 1}$ is a predictable stepsize sequence. Formally, we define the filtration $\mathcal{F}_{t}:=\sigma(x_{1},g_{1},\ldots,x_{t},g_{t})$, for $t\geq 1$, and assume that $\eta_{t}$ is $\mathcal{F}_{t-1}$ measurable. Because $x_{1}\in\mathcal{X}$ and each update projects back onto $\mathcal{X}$, we have $x_{t}\in\mathcal{X}$ for all $t\geq 1$. In particular, all iterates lie in a bounded region around $x^{\star}$:
$\|x_{t}-x^{\star}\|\;\leq\;R_{x}:=\textrm{diam}(\mathcal{X})<\infty.$
The stochastic gradients $g_{t}$ are assumed to satisfy the following three conditions.
Assumption 1.1 (Stochastic gradients).
@@
    Conditional second moment bound: $\mathbb{E}[\|g_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq G^{2}$ almost surely.
  • (iii)
@@
    $\mathbb{E}\!\left[\exp\!\big(\lambda\langle u,\xi_{t}\rangle\big)\,\Big|\,\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\sigma^{2}\|u\|^{2}}{2}\right)\qquad\forall u\in\mathbb{R}^{d},\ \forall\lambda\in\mathbb{R}.$ (2)
Assumption 1.1(i) is the standard unbiasedness condition used in stochastic approximation. Assumption 1.1(ii) requires a uniform bound on the conditional second moment of the stochastic gradients, which is classical in nonasymptotic analyses of SGD and ensures that the quadratic term $\|g_{t}\|^{2}$ remains controlled. Finally, Assumption 1.1(iii) postulates that the gradient noise $\xi_{t}\coloneqq g_{t}-\nabla f(x_{t})$ is conditionally sub-Gaussian given the past. This condition provides a quantitative tail bound on the one-step fluctuations of the stochastic gradients and is natural in time-uniform concentration arguments. Moreover, it is satisfied by a broad class of noise models, including bounded, Gaussian, and many light-tailed stochastic gradients frequently encountered in practice. We operate in the standard stochastic approximation regime:
Assumption 1.2 (Stepsizes).
The predictable stepsizes $\{\eta_{t}\}_{t\geq 1}$ satisfy
$\sum_{t=1}^{\infty}\eta_{t}=\infty,\qquad\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty,\qquad 0<\eta_{t}<\infty.$ (3)
These conditions ensure that the algorithm continues making progress in early iterations while the cumulative noise remains controlled. It is convenient to introduce two standard cumulative quantities from stochastic approximation,
@@
which will appear naturally in our analysis. To measure optimization progress, we focus on the weighted average suboptimality,
$\bar{F}_{t}:=\frac{\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)}{S_{t}}.$ (4)
This statistic is classical in stochastic approximation theory and has a direct optimization interpretation. Indeed, defining the corresponding weighted average iterate
$\bar{x}_{t}:=\frac{1}{S_{t}}\sum_{s=1}^{t}\eta_{s}x_{s},$ (5)
@@
Thus $\bar{F}_{t}$ simultaneously controls the performance of the averaged iterate $\bar{x}_{t}$ and serves as the natural target for our anytime-valid analysis.
@@ 1.2 Contributions
Our goal is to provide anytime-valid, data-dependent guarantees for the progress of projected stochastic gradient descent. Specifically, given the stochastic gradients $\{g_{t}\}$, the predictable stepsizes $\{\eta_{t}\}$, and the observed iterates $\{x_{t}\}$, we construct an upper confidence sequence for the weighted average suboptimality $\bar{F}_{t}$ in (4), which holds uniformly over all times $t\geq 1$. This confidence sequence depends on trajectory-dependent quantities: the stochastic gradient norms, the stepsizes, and the squared-distance process $\|x_{t}-x^{\star}\|^{2}$. While the latter is not directly observable, it is uniformly bounded by $R_{x}^{2}$ due to compactness of $\mathcal{X}$, which immediately yields a fully implementable bound depending only on the realized gradient norms, the stepsizes, and $R_{x}$. To state our contributions precisely, we recall the formal notion of an anytime-valid upper confidence sequence.
@@
Let $\{\bar{F}_{t}\}_{t\geq 1}$ be as in (4). A sequence of random variables $\{U_{t}(\alpha)\}_{t\geq 1}$ is an anytime-valid upper confidence sequence for $\{\bar{F}_{t}\}$ if
$\mathbb{P}\!\left(\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\right)\geq 1-\alpha.$ (6)
These uniform-in-time bounds enable a natural data-dependent stopping rule. Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,1)$, we define the stopping time
$\tau_{\varepsilon}:=\inf\left\{t\geq 1:U_{t}(\alpha)\leq\varepsilon\right\},$ (7)
which returns the first iterate that is certified to be $\varepsilon$-optimal.
Main contributions.
Our main contributions can be summarized as follows.
@@
    Anytime-valid confidence sequence. Fix $\alpha\in(0,\tfrac{2}{\mathrm{e}})$. Theorem 2.1 provides an upper confidence sequence $\{U_{t}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ satisfying
    $\mathbb{P}\!\left(\forall t\geq 1:\ \bar{F}_{t}\leq U_{t}(\alpha)\right)\geq 1-\alpha.$
    The bound $U_{t}(\alpha)$ is trajectory-adaptive and remains valid uniformly over time, including under adaptive monitoring and data-dependent choices of stepsizes. Moreover, a fully observable version of the confidence sequence, suitable for online implementation, is given in Corollary 2.4. Finally, under standard stochastic approximation stepsizes, $U_{t}(\alpha)$ decays at the rate $1/S_{t}$ almost surely; see Proposition 2.2 and Remark 2.3.
@@
    Anytime-valid $\varepsilon$-optimal stopping rule. Fix $\varepsilon>0$ and $\alpha\in(0,\tfrac{2}{\mathrm{e}})$. Theorem 3.1 proves that the stopping time $\tau_{\varepsilon}$ defined in (7) is almost surely finite, and the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies
    $\mathbb{P}\!\left(f(\bar{x}_{\tau_{\varepsilon}})-f(x^{\star})\leq\varepsilon\right)\geq 1-\alpha.$
    Thus projected SGD, when equipped with the confidence sequence from (i), becomes an anytime-valid optimization procedure with a rigorous finite-time certificate of $\varepsilon$-optimality. Explicit bounds on the stopping-time complexity for standard stepsize schedules are given in Propositions 3.4 and 3.5.
@@ 1.3 Related Work
Conceptually, our work lies at the intersection of two developments: (i) time-uniform inference, which provides a framework for hypothesis tests and confidence intervals that yields valid control over errors uniformly in time, and (ii) stochastic approximation, which delivers nonasymptotic convergence guarantees for stochastic gradient methods in convex optimization. The present paper combines these perspectives by adapting nonnegative supermartingale techniques from time-uniform inference to the projected SGD setting. This yields a fully data-dependent, anytime-valid upper confidence sequence for suboptimality, together with a stopping rule that is provably $\varepsilon$-optimal with high probability and almost surely finite under standard stochastic approximation conditions. In what follows, we position our contributions within the broader literature on stochastic approximation, averaging and weighted iterates, optimization stopping criteria, time-uniform concentration, and confidence-sequence methodology.
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [robbins1951stochastic] established almost sure convergence of iterative schemes under diminishing stepsizes, while [kushner2003stochastic] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [polyak1992acceleration]. This asymptotic perspective was sharpened in [duchi2021asymptotic], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [davis2024asymptotic]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established high-probability bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [nemirovski2009robust, lan2012optimal]. Related work also studied statistical validation of mirror-descent–type stochastic approximation methods, providing confidence bounds for the optimal objective value at a fixed iteration horizon rather than under adaptive or continuously monitored stopping rules [lan2012validation]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias–variance tradeoffs and finite-time rates for averaged iterates [moulines2011non, bach2013non, dieuleveut2016nonparametric]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [rakhlin2012making, shamir2013stochastic, harvey2019tight, harvey2019simple, jain2021making]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [agarwal2012information]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, including heavy-tailed settings handled via gradient clipping and related robustification techniques [gorbunov2020stochastic, cutkosky2021high], as well as general boosting and robust-estimation wrappers that convert low-probability guarantees into high-confidence ones [davis2021low]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper. Viewed differently, the present work does not seek to refine convergence rates for SGD, but rather to introduce a complementary, time-uniform notion of statistical validity for stochastic optimization algorithms, in which stopping times are treated as random, data-dependent objects subject to formal inference.

### v4

# Stopping Rules for SGD via Anytime-Valid Confidence Sequences

@@ Abstract
Deciding when to stop stochastic gradient descent (SGD) has long remained unresolved in a statistically rigorous sense. While SGD is routinely monitored as it runs, the classical theory of SGD provides guarantees only at pre-specified iteration horizons and offers no valid way to decide, based on the observed trajectory, when further computation is justified. We address this gap by developing anytime-valid confidence sequences for stochastic gradient methods, which remain valid under continuous monitoring and directly induce statistically valid, trajectory-dependent stopping rules: stop as soon as the current upper confidence bound on an appropriate performance measure falls below a user-specified tolerance. The confidence sequences are constructed using nonnegative supermartingales, are time-uniform, and depend only on observable quantities along the SGD trajectory, without requiring prior knowledge of the optimization horizon. In convex optimization, this yields anytime-valid certificates for weighted suboptimality of projected SGD under general stepsize schedules, without assuming smoothness or strong convexity. In nonconvex optimization, it yields time-uniform certificates for weighted first-order stationarity under smoothness assumptions. We further characterize the stopping-time complexity of the resulting stopping rules under standard stepsize schedules. To the best of our knowledge, this is the first framework that provides statistically valid, time-uniform stopping rules for SGD across both convex and nonconvex settings based solely on its observed trajectory.
@@ 1 Introduction
Deciding when to stop an optimization algorithm is often treated as a routine design choice. One selects a number of iterations, runs the algorithm, and accepts the resulting iterate. Yet this convention conceals a basic tension. The chosen horizon is necessarily arbitrary: had the algorithm converged more quickly, computation could have been saved; had progress been slower, additional iterations might have led to a meaningfully better solution. Standard optimization theory offers little guidance for resolving this tension while the algorithm is actually running.
This issue is particularly acute for stochastic gradient descent (SGD), which in practice is almost always monitored as it unfolds. Existing analyses of SGD characterize its behavior by answering a forward-looking question: if the algorithm is run for a prescribed number of iterations, what guarantees can be made about optimality or stationarity? In practice, however, users face the inverse problem. At a given iteration, after observing the realized trajectory of the algorithm, one would like to know whether continuing to iterate is statistically justified. This inversion of perspective lies at the heart of practical optimization, where computational budgets are finite and decisions must be made online.
A natural response is to monitor the algorithm’s progress as it unfolds and to stop once some criterion appears satisfactory. However, this adaptive monitoring fundamentally alters the statistical setting. Guarantees derived for fixed iteration horizons do not remain valid when the algorithm is inspected repeatedly and stopping decisions are based on the observed trajectory. Classical confidence bounds implicitly assume that the stopping rule is fixed in advance; once this assumption is violated, repeated checking induces multiple testing effects that invalidate standard statistical conclusions. As a result, widely used stopping heuristics, including fixed budgets, visually detected plateaus, or ad hoc thresholds, provide no quantifiable protection against premature or overly conservative termination.
The consequences are not merely theoretical. In large-scale learning, reinforcement learning, and scientific computing, the cost of unnecessary computation can be substantial, while stopping too early may leave significant performance gains unrealized. What is missing is a mechanism that translates the observed trajectory of SGD into statistically reliable, real-time evidence about progress, without requiring the stopping time to be specified in advance. This raises the central question of this paper:
How close is the current iterate to optimality (or stationarity)?
In this work, we address this gap by formulating stopping for SGD as a problem of sequential statistical inference. Our approach is based on anytime-valid confidence sequences, which are designed to remain valid under continuous monitoring and arbitrary data-dependent stopping. By adapting this framework to the dynamics of stochastic gradient methods, we obtain time-uniform, data-dependent certificates of performance that can be inspected at every iteration without invalidating their guarantees. These certificates directly give rise to simple, statistically valid stopping rules that depend only on the observed trajectory of the algorithm.
This framework applies to both convex and nonconvex optimization. In convex settings, it yields anytime-valid certificates for suboptimality that allow one to stop SGD with a guaranteed level of optimality. In nonconvex settings, where global optimality is unattainable in general, it yields time-uniform certificates for first-order stationarity. In both cases, the resulting stopping rules are compatible with standard stepsize schedules and preserve their validity under arbitrary adaptive monitoring.
@@ 1.1 Problem formulation
We consider stochastic optimization problems of the form
$\min_{x\in\mathcal{X}}f(x),$
where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a closed convex set and $f:\mathcal{X}\to\mathbb{R}$ is a (possibly nonconvex) objective function. Throughout, we study stochastic gradient descent with projection onto $\mathcal{X}$, defined recursively by
$x_{t+1}=\Pi_{\mathcal{X}}(x_{t}-\eta_{t}g_{t})\qquad\forall t\geq 1,$ (1)
where $\Pi_{\mathcal{X}}$ denotes Euclidean projection onto $\mathcal{X}$, $\{g_{t}\}_{t\geq 1}$ are stochastic gradients, and $\{\eta_{t}\}_{t\geq 1}$ is a stepsize sequence. We define the natural filtration $\mathcal{F}_{t}:=\sigma(x_{1},g_{1},\ldots,x_{t},g_{t})$, for $t\geq 1$, and assume that the stepsizes are predictable with respect to this filtration.
Assumption 1.1 (Stepsizes).
The stepsizes $\{\eta_{t}\}_{t\geq 1}$ are $\mathcal{F}_{t-1}$-measurable for all $t\geq 1$.
Assumption 1.1 allows for arbitrary data-dependent stepsize rules, provided they depend only on the past trajectory. In particular, it covers standard constant stepsizes, diminishing schedules such as $\eta_{t}=\eta_{0}\,t^{-1/2}$, classical stochastic approximation stepsizes satisfying $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$, as well as adaptive stepsizes that are functions of the observed iterates and gradients (e.g., AdaGrad-type schedules and related adaptive methods).
The stochastic gradients are assumed to satisfy the following conditions, which are common to both the convex and nonconvex settings.
Assumption 1.2 (Stochastic gradients I).
@@
    $\mathbb{E}\!\left[\exp\!\big(\lambda\langle u,\xi_{t}\rangle\big)\,\Big|\,\mathcal{F}_{t-1}\right]\leq\exp\!\left(\frac{\lambda^{2}\sigma^{2}\|u\|^{2}}{2}\right)\qquad\text{almost surely }\forall u\in\mathbb{R}^{d},\forall\lambda\in\mathbb{R}.$ (2)
Assumption 1.2(i) is the standard unbiasedness condition used in stochastic approximation. Moreover, Assumption 1.2(ii) postulates that the gradient noise is conditionally sub-Gaussian given the past. This condition provides quantitative tail control on the one-step fluctuations of the stochastic gradients and is natural in time-uniform concentration arguments; it is satisfied by bounded, Gaussian, and many light-tailed noise models encountered in practice.
For convenience, we introduce the cumulative quantities
@@
which will appear repeatedly in the analysis. We now distinguish between the convex and nonconvex settings, which differ only in the choice of performance measure and in the additional assumptions imposed on the objective in the nonconvex case.
Convex case. In the convex setting, we assume that $f$ is convex and that the set $\mathcal{X}$ is compact. This assumption is used only to guarantee a uniform bound on the distance $\|x_{t}-x^{\star}\|$ along the SGD trajectory; see Remark 2.4 for a discussion of relaxed conditions. Let
$x^{\star}\in\arg\min_{x\in\mathcal{X}}f(x)$
denote an arbitrary minimizer. Since all iterates satisfy $x_{t}\in\mathcal{X}$, they remain in a bounded neighborhood of $x^{\star}$:
$\|x_{t}-x^{\star}\|\leq R_{x}:=\operatorname{diam}(\mathcal{X})<\infty.$
To measure optimization progress, we focus on the weighted average suboptimality
$\bar{F}_{t}:=\frac{1}{S_{t}}\sum_{s=1}^{t}\eta_{s}\bigl(f(x_{s})-f(x^{\star})\bigr).$ (3)
Defining the corresponding weighted average iterate
$\bar{x}_{t}:=\frac{1}{S_{t}}\sum_{s=1}^{t}\eta_{s}x_{s},$
@@
Thus $\bar{F}_{t}$ simultaneously controls the performance of the averaged iterate and serves as the natural target for our anytime-valid analysis in the convex case.
Nonconvex case. In the nonconvex setting, we take $\mathcal{X}=\mathbb{R}^{d}$ and impose no convexity or coercivity assumptions on the objective. We assume that $f:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is differentiable. The nonnegativity assumption is made without loss of generality by shifting the objective if necessary. We do, however, require a standard smoothness condition on the objective.
Assumption 1.3 (Smoothness).
The function $f$ is $L$-smooth, that is,
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|,\qquad\forall x,y\in\mathbb{R}^{d}.$
In the absence of convexity, optimality gaps are no longer an appropriate performance measure. Instead, we quantify progress through a weighted stationarity measure,
$\bar{G}_{t}:=\frac{1}{S_{t}}\sum_{s=1}^{t}\eta_{s}\|\nabla f(x_{s})\|^{2}.$ (4)
This quantity is standard in nonconvex stochastic optimization and captures the extent to which the algorithm has approached a stationary regime. Moreover, since the stepsizes are nonnegative, $\bar{G}_{t}$ upper bounds the smallest squared gradient norm encountered along the trajectory:
$\min_{1\leq s\leq t}\|\nabla f(x_{s})\|^{2}\leq\bar{G}_{t}.$
Our goal in the nonconvex setting is to construct time-uniform, data-dependent upper confidence bounds on $\bar{G}_{t}$ that remain valid under continuous monitoring and arbitrary data-dependent stopping.
@@ 1.2 Contributions
Our goal is to provide anytime-valid, data-dependent guarantees for the progress of stochastic gradient descent across both convex and nonconvex optimization settings. Specifically, given the stochastic gradients $\{g_{t}\}$, the predictable stepsizes $\{\eta_{t}\}$, and the observed iterates $\{x_{t}\}$, we construct time-uniform upper confidence bounds for a suitable performance measure associated with the SGD trajectory. In the convex case, this measure is the weighted average suboptimality $\bar{F}_{t}$ defined in (3), while in the nonconvex case it is the weighted average stationarity measure $\bar{G}_{t}$ defined in (4). In both settings, the resulting bounds remain valid simultaneously for all times $t\geq 1$, including under continuous monitoring and arbitrary data-dependent stopping. To state our contributions precisely, we recall the formal notion of an anytime-valid upper confidence sequence.
@@
Let $\{\bar{H}_{t}\}_{t\geq 1}$ denote a nonnegative performance process, such as $\bar{F}_{t}$ in the convex case or $\bar{G}_{t}$ in the nonconvex case. A sequence of random variables $\{U_{t}(\alpha)\}_{t\geq 1}$ is an anytime-valid upper confidence sequence for $\{\bar{H}_{t}\}$ if
$\mathbb{P}\!\left(\forall t\geq 1:\ \bar{H}_{t}\leq U_{t}(\alpha)\right)\geq 1-\alpha.$
Such time-uniform bounds naturally induce statistically valid stopping rules. Given a target accuracy $\varepsilon>0$ and confidence level $\alpha\in(0,1)$, we define the stopping time
$\tau_{\varepsilon}:=\inf\left\{t\geq 1:U_{t}(\alpha)\leq\varepsilon\right\},$
which returns the first iterate that is certified to have reached the desired performance level. In the convex case, this corresponds to $\varepsilon$-optimality, while in the nonconvex case it corresponds to $\varepsilon$-stationarity (up to the appropriate scaling of the performance measure).
Main contributions. Our main contributions can be summarized as follows.
@@
    Anytime-valid confidence sequences. Under the assumptions of Section 1.1, we construct anytime-valid upper confidence sequences $\{U_{t}(\alpha)\}_{t\geq 1}$, in the sense of Definition 1.4, for the weighted average suboptimality $\bar{F}_{t}$ in the convex setting and for the weighted average stationarity measure $\bar{G}_{t}$ in the nonconvex setting. The resulting bounds are trajectory-adaptive, depend only on quantities generated along the SGD path, and remain valid uniformly over time, including under continuous monitoring and data-dependent choices of stepsizes. In the convex case, compactness of the feasible set yields a fully observable confidence sequence depending only on realized stochastic gradient norms, the stepsizes, and the domain diameter. In the nonconvex case, the confidence sequence controls $\bar{G}_{t}$ and becomes fully observable under additional boundedness assumptions on $\|\nabla f(x_{t})\|$. Finally, we characterize the rates of decay of these anytime-valid certificates under common stepsize choices, including (square-summable) stochastic approximation stepsizes and the canonical $\eta_{t}=\eta_{0}\,t^{-1/2}$ schedule, obtaining explicit bounds in expectation and almost surely along the realized SGD trajectory.
@@
    Statistically valid stopping rules. The anytime-valid confidence sequences in (i) directly induce trajectory-dependent stopping rules that certify $\varepsilon$-optimality in the convex case and $\varepsilon$-stationarity in the nonconvex case with confidence $1-\alpha$, without requiring a predetermined optimization horizon. These guarantees hold for arbitrary predictable stepsize sequences, including adaptive stepsizes that are functions of the observed iterates and stochastic gradients. To provide concrete insight into the resulting stopping behavior, we analyze the induced certified stopping rules by deriving expected stopping-time bounds (i.e., bounds on $\mathbb{E}[\tau_{\varepsilon}]$) under commonly used stepsize schedules, including $\eta_{t}=\eta_{0}t^{-1/2}$ and more general stochastic approximation–type schedules.
@@ 1.3 Related work
Conceptually, this work sits at the intersection of two lines of research. The first concerns time-uniform statistical inference, which develops hypothesis tests and confidence bounds that remain valid under continuous monitoring and arbitrary data-dependent stopping. The second concerns stochastic approximation, which provides convergence guarantees for stochastic gradient methods in both convex and nonconvex optimization. The present paper brings these perspectives together by adapting nonnegative supermartingale techniques from time-uniform inference to the dynamics of stochastic gradient descent. Below, we position our contributions within the broader literature on time-uniform statistical inference and stochastic optimization.